# Option Pricing with Time-Varying Volatility Risk Aversion

### v3

@@ Abstract
We introduce a pricing kernel with time-varying volatility risk aversion that can explain the observed time variation in the shape of the pricing kernel. Dynamic volatility risk aversion, combined with the Heston-Nandi GARCH model, leads to a convenient option pricing model, denoted DHNG. The variance risk ratio emerges as a fundamental variable, and we show that it is closely related to economic fundamentals and common measures of sentiment and uncertainty. DHNG yields a closed-form pricing formula for the VIX, and we propose a novel approximation method that provides analytical expressions for option prices. We estimate the model using S&P 500 returns, the VIX, and option prices, and find that dynamic volatility risk aversion leads to a substantial reduction in VIX and option pricing errors.
1 Introduction
@@
The pricing kernel is the ratio of risk-neutral state probabilities to physical state probabilities, which is a fundamental concept in economics with roots in Arrow-Debreu securities, see Arrow and Debreu, (1954). A cornerstone of asset pricing theory is the Lucas, (1978) tree model, which implies a pricing kernel that is monotonically decreasing in aggregate wealth, see Hansen and Renault, (2010). The pricing kernel can be estimated from the underlying asset and derivative prices, see e.g. Aït-Sahalia and Lo, (2000), Jackwerth, (2000) and Rosenberg and Engle, (2002), and this literature has uncovered a number of stylized facts that are known as pricing kernel puzzles. One puzzle is that the empirical pricing kernel is upward-sloping in some regions when plotted against the market return, which is at odds with the Lucas tree model and several modern option pricing models.
@@
The discrete-time Heston-Nandi GARCH (HNG) model, see Heston and Nandi, (2000), conveniently yields closed-form expressions for option prices while allowing for time-varying volatility, but it also implies a monotonic pricing kernel. Christoffersen et al., (2013) generalized the HNG model to include a variance-dependent pricing kernel, which results in a non-monotonic pricing kernel. Their model is also tractable, yields closed-form expressions for option prices, and has become a benchmark in the option pricing literature. The shape of the pricing kernel in Christoffersen et al., (2013) is defined by the variance risk premium (VRP), with a negative VRP implying a U-shaped pricing kernel. This is the shape they find in their empirical analysis, see also Cuesdeanu and Jackwerth, (2018).
@@
While the framework in Christoffersen et al., (2013) can resolve one pricing puzzle, it conflicts with another stylized fact: the shape of the pricing kernel varies over time. Typically, the pricing kernel exhibits a U-shape when estimated over long sample periods, but it has distinctly different shapes during certain periods, such as the years leading up to the global financial crisis. The shape of the pricing kernel can be expressed in terms of the probability weighting function. If individuals overweight low-probability events and underweight high-probability events, the probability weighting function will have an inverse S-shape, which corresponds to a U-shaped pricing kernel. Polkovnichenko and Zhao, (2013) estimated the probability weighting function non-parametrically and found it to be time-varying. Their estimate has an inverse S-shape during most of their sample period but a regular S-shape during the years 2004–2006. The latter is consistent with the negative risk aversion function documented by Aït-Sahalia and Lo, (2000) and Jackwerth, (2000). Similar results were obtained by Chabi-Yo and Song, (2013), and the inverted U-shape was also documented with DAX 30 options during the same years, 2004–2006, see Grith et al., (2017). Time variation was also found by Kiesel and Rahe, (2017) and by Beare and Schmidt, (2016). The latter developed a formal statistical test for monotonicity and used it to identify days with unusual shapes. Interestingly, Christoffersen et al., (2013, figure 3) also contains evidence of a pricing kernel whose shape varies over time. For most calendar years, their estimated empirical pricing kernel is U-shaped; however, in some calendar years, it takes an inverted U-shape. [^1] The authors do not comment on this observation, but they estimate a more flexible structure, which is not based on a pricing kernel with a fixed shape, see Christoffersen et al., (2013, figure 6). The inverted U-shape can be induced by (unmodeled) parameter instability, such as structural breaks, see Sichert, (2022), and Markov switching dynamics, see Tong et al., (2022).
@@
In this paper, we make five main contributions. First, we propose a novel framework for derivatives pricing based on a pricing kernel with time-varying volatility risk aversion. When this pricing kernel is combined with the GARCH model of Heston and Nandi, (2000), the risk-neutral conditional variance divided by the physical conditional variance, $\eta_{t}=h_{t+1}^{\ast}/h_{t+1}$, emerges as a fundamental quantity. We refer to $\eta_{t}$ as the variance risk ratio (VRR), which is closely related to the variance risk premium. The VRR governs time variation in the pricing kernel and is functionally linked to the volatility risk aversion parameter that defines the curvature of the pricing kernel. We demonstrate that this structure can generate the observed time variation in the empirical pricing kernel, including its various shapes. We label this model as DHNG, where “D” references the dynamic shape of the pricing kernel and related quantities. DHNG nests the model by Christoffersen et al., (2013), which we denote as CHNG, as it corresponds to the special case where $\eta_{t}$ is constant. The HNG model is nested as the special case where $\eta_{t}=1$, for all $t$.
@@
Second, we derive a closed-form CBOE VIX pricing formula for the DHNG model and analytical option pricing formulas, where the latter entails a methodological contribution. When volatility risk aversion follows a stochastic process, the moment generating function (MGF) for cumulative returns does not have an affine form. It is nevertheless simple to obtain a closed-form expression for the VIX, whereas deriving option pricing formulas is more challenging. We propose a novel approximation method, which is not specific to our model, that leads to an analytical option pricing formula. The underlying idea is to identify an auxiliary MGF for a simpler, related problem and deduce option prices from a suitable Taylor expansion of this auxiliary MGF. We find the analytical approximation method to be highly accurate in an empirically relevant simulation design.
@@
The third contribution is a practical implementation of the theoretical framework. We connect the theoretical model for $\eta_{t}$ with observed data using an observation-driven model, which is a natural choice because the GARCH model under the physical measure has this structure. Specifically, we propose a score-driven model, see Creal et al., (2013), where the first-order conditions of the log-likelihood function define the innovations in the dynamic model for $\eta_{t}$. This approach is intuitive because the volatility risk ratio, $\eta_{t}$, is updated with the objective of minimizing pricing errors. Conveniently, this design adds robustness to model misspecification. A key advantage of observation-driven models is their simplicity in estimation, which is particularly useful in applications involving a long sample period and a large panel of option prices.
@@
The fourth contribution is an empirical application using 32 years of daily S&P 500 returns, the CBOE VIX, and a large panel of option prices. The estimation incorporates information from both the physical and risk-neutral probability measure, as advocated by Chernov and Ghysels, (2000). We find that the dynamic pricing kernel with time-varying volatility risk aversion typically reduces pricing errors by 50% or more, as measured by the root mean square error. This reduction is achieved for both VIX prices and option prices, is observed both in-sample and out-of-sample. The estimated time variation in the shape of the pricing kernel is consistent with existing studies on this topic. The estimated DHNG model generally exhibits a U-shaped pricing kernel; however, an inverted U-shape is observed during the years 2004 to 2007, as well as in periods around 1993 and 2017.
@@
Fifth, we connect the VRR with several key asset pricing variables. Because the VRR emerges as a fundamental variable in this framework, it is interesting to investigate its relationships with other variables used in this literature. Several variables have been tied to pricing kernel puzzles. For instance, investors with heterogeneous beliefs and disagreements about the physical distribution can result in a U-shaped pricing kernel. The U-shape arises as a consequence of the aggregation of investors who take opposite positions in the market due to differing beliefs about expected returns. Time variation in the level of disagreement translates to time variation in the pricing kernel, see e.g. Shefrin, (2001, 2008), Bakshi and Madan, (2008), Bakshi et al., (2010). A similar argument can be made about sentiment and uncertainty among investors, see e.g. Han, (2008), Polkovnichenko and Zhao, (2013), Baker et al., (2016), and Bollerslev et al., (2018). Bali and Zhou, (2016) argue that market uncertainty can be approximated by the variance risk premium, see Carr and Wu, (2008). The VRR is obviously related to the variance risk premium, which has been shown to predict stock returns, see Bollerslev et al., (2009). A straightforward explanation is that preferences are state-dependent with market uncertainty as a possible state variable, see e.g. Grith et al., (2017). We connect our results with these studies by showing that the monthly average VRR is closely related to commonly used measures of sentiment, uncertainty, and disagreement, as well as other economic variables that are commonly used in this literature.
@@
Our paper is related to, and builds on, a large body of literature. This includes the literature that sought to explain the pricing kernel puzzles by augmenting existing models with additional state variables, see e.g. Chabi-Yo et al., (2008), Chabi-Yo, (2012), Brown and Jackwerth, (2012) and Song and Xiu, (2016). The VRR can be interpreted as an additional state variable that emerges naturally from our model structure. Our paper is also related to Barone-Adesi et al., (2008), who estimated a GJR-GARCH model for returns under the physical measure. They did not specify a pricing kernel but simply assumed that the model under the risk-neutral measure is also a GJR-GARCH model, whose time-varying parameters were calibrated to minimize pricing errors in daily options. We also seek to minimize pricing errors, but our approach is fundamentally different. Our starting point is a pricing kernel and a model for the physical probability measure, $\mathbb{P}$, and the two imply the model under the risk-neutral probability measure, $\mathbb{Q}$. In their empirical comparison, Christoffersen et al., (2013) explored the same ideas to Barone-Adesi et al., (2008) by estimating two separate HNG models under both $\mathbb{P}$ and $\mathbb{Q}$. They referred to this structure as an ad-hoc model, because it is not deduced from their pricing kernel. We show that the pricing kernel with a time-varying VRR is not coherent with HNG models under both $\mathbb{P}$ and $\mathbb{Q}$.
@@
Our paper is also related to the literature on time-varying risk aversion because it is connected to the local shape of the pricing kernel. A seminal paper on time-varying risk aversion is Campbell and Cochrane, (1999), and Li, (2007) showed how time-varying risk aversion relates to a time-varying risk premium in his framework. Moreover, González et al., (2018) found that time-varying risk aversion is one of the main determinants of stock market betas. A recent paper by Bekaert et al., (2022) introduced a no-arbitrage asset pricing model with time-varying risk aversion for the pricing of equities and corporate bonds, where the risk aversion is driven by the uncertainty shocks.
@@
The rest of the paper is organized as follows. We present the theoretical model in Section 2 and derive closed-form pricing formula for the VIX and an option pricing formula, based on the new approximation method in Section 3. Section 4 introduces a dynamic model for the variance risk ratio using a score-driven approach. An empirical application with 32 years of S&P 500 returns, VIX, and option prices is presented in Section 5. We show that the variance risk ratio relates to well-known measures of sentiment, disagreement, and uncertainty in Section 6. A summary and concluding remarks are presented in Section 7. All proofs are provided in the Appendix A, and supplementary empirical results are in Appendix B.
@@ 2 The Model
The observed variables include the daily returns, $R_{t}\equiv\log\left(S_{t}/S_{t-1}\right)$, where $S_{t}$ is the underlying asset price, and a vector of derivative prices, denoted $X_{t}$. We will work with two filtrations, $\mathcal{F}_{t}=\sigma(\{R_{j},X_{j}\},j\leq t)$ and $\mathcal{G}_{t}=\sigma(\{R_{j},X_{j-1}\},j\leq t)$, where the latter arises naturally from a factorization of the joint likelihood function. Clearly, $\mathcal{G}_{t}\subset\mathcal{F}_{t}$ for all $t$. We model returns in the physical measure $\mathbb{P}$ using the classical Heston-Nandi GARCH model, which has a convenient structure for derivatives pricing. This model is given by
@@
where the return shock, $z_{t+1}$, is assumed to be independent and identically distributed (iid) with a standard normal distribution, $N(0,1)$; $r$ is the risk free rate; $\lambda$ is the equity risk premium because the expected return is given by $\mathbb{E}\left(\exp\left(R_{t+1}\right)|\mathcal{F}_{t}\right)=\exp\left(r+\lambda h_{t+1}\right)$; and $h_{t+1}=\mathrm{var}(R_{t+1}|\mathcal{F}_{t})$ is the daily conditional variance, see Heston and Nandi, (2000). The HNG model captures time variation in the conditional variance, as is the case for ARCH and GARCH models, see Engle, (1982) and Bollerslev, (1986). However, the HNG model also allows for dependence between returns and volatility (leverage). It is straightforward to verify that $\mathrm{cov}_{t-1}(R_{t},h_{t+1})=-2\gamma\alpha h_{t}$, which shows that the magnitude of the leverage effect is defined by $\gamma$. A leverage effect is required to generate the empirically important volatility smirk in option prices. The dynamic structure of the HNG model is carefully crafted to yield a closed-form solution for option valuation, and Heston and Nandi, (2000) showed that the continuous limit of the HNG model (as the time interval between observations shrinks to zero) yields a variance process, $h_{t}$, that converges weakly to a continuous-time square-root variance process, see Feller, (1951), Cox et al., (1985), and Heston, (1993).

### v4

# Peter Reinhard Hansena and Chen Tongb,c

@@ Abstract
We introduce a pricing kernel with time-varying volatility risk aversion to explain observed time variations in the shape of the pricing kernel. When combined with the Heston-Nandi GARCH model, this framework yields a tractable option pricing model in which the variance risk ratio (VRR) emerges as a key variable. We show that the VRR is closely linked to economic fundamentals, as well as sentiment and uncertainty measures. A novel approximation method provides analytical option pricing formulas, and we demonstrate substantial reductions in pricing errors through an empirical application to the S&P 500 index, the CBOE VIX, and option prices.
Introduction
@@
The pricing kernel, which represents the ratio of risk-neutral to physical state probabilities, is a fundamental concept in economics with roots in Arrow-Debreu securities, see Arrow and Debreu, (1954). A cornerstone of asset pricing theory, the Lucas, (1978) tree model, implies a monotonically decreasing pricing kernel with respect to aggregate wealth, see Hansen and Renault, (2010). However, empirical studies consistently find the pricing kernel to be non-monotonic and time-varying, giving rise to the so-called pricing kernel puzzles, see e.g. Aït-Sahalia and Lo, (2000), Jackwerth, (2000), and Rosenberg and Engle, (2002).
In this paper, we propose a pricing kernel with time-varying volatility risk aversion to resolve these pricing kernel puzzles. Building on Heston and Nandi, (2000) and Christoffersen et al., (2013), we develop a novel framework for derivatives pricing, in which the variance risk ratio (VRR) emerges as the key quantity. The VRR, defined as the ratio of risk-neutral to physical conditional variance, characterizes both the shape and time variation of the pricing kernel and it is closely related to the variance risk premium by Carr and Wu, (2008). The dynamic pricing kernel introduces challenges for closed-form option pricing formulas, which we address through a novel analytical approximation method applicable to non-affine models. Our empirical implementation of the dynamic structure for VRR is based on the score-driven framework of Creal et al., (2013), which intuitively updates time-varying quantities in response to first-order conditions. In an empirical application to the S&P 500 index, the CBOE VIX, and option prices, we show that the new model substantially reduces pricing errors, both in-sample and out-of-sample.
@@
The discrete-time Heston-Nandi GARCH (HNG) model, developed by Heston and Nandi, (2000), conveniently yields closed-form expressions for option prices while allowing for time-varying volatility. However, it implies a monotonic pricing kernel. Christoffersen et al., (2013) extended the HNG model by introducing a variance-dependent pricing kernel, resulting in a non-monotonic shape. Their model remains tractable, retains closed-form option pricing expressions, and has become a benchmark in the option pricing literature. The shape of the pricing kernel in their framework is governed by the variance risk premium (VRP), where a negative VRP leads to a U-shaped pricing kernel, a pattern confirmed in their empirical analysis, see also Cuesdeanu and Jackwerth, (2018).
@@
While Christoffersen et al., (2013) explain non-monotonicity, their framework cannot account for the time variation observed in the pricing kernel. When estimated over long sample periods, the pricing kernel is typically U-shaped, yet the shape differs distinctly in certain periods, such as the years leading up to the global financial crisis. The shape of the pricing kernel can be expressed in terms of the probability weighting function. If individuals overweight low-probability events and underweight high-probability events, the probability weighting function will have an inverse S-shape, which corresponds to a U-shaped pricing kernel. Polkovnichenko and Zhao, (2013) estimated the probability weighting function non-parametrically and found it to be time-varying. Their estimate has an inverse S-shape during most of their sample periods but a regular S-shape during the years 2004–2006. Similar results were obtained by Chabi-Yo and Song, (2013), and an inverted U-shape was also documented with DAX 30 options during the same years, 2004–2006, as reported by Grith et al., (2017). Further evidence comes from Kiesel and Rahe, (2017) and Beare and Schmidt, (2016). Interestingly, Christoffersen et al., (2013, figure 3) also contains evidence of time variations in empirical pricing kernels. For most calendar years, their estimated empirical pricing kernel is U-shaped; however, in some calendar years, it takes an inverted U-shape. [^1] The authors do not comment on this observation, but they estimate a more flexible structure, which is not based on a pricing kernel with a fixed shape, see Christoffersen et al., (2013, figure 6).
@@
The derivative pricing model proposed in this paper combines a pricing kernel with time-varying volatility risk aversion and the GARCH model of Heston and Nandi, (2000). The VRR, defined as the ratio of risk-neutral to physical conditional variance, $\eta_{t}=h_{t+1}^{\ast}/h_{t+1}$, governs the time variation in the pricing kernel and is functionally linked to the volatility risk aversion parameter, which captures the curvature of the pricing kernel. We demonstrate that this structure can generate the observed time variation in the empirical pricing kernel, including its different shapes. Our new model nests Christoffersen et al., (2013) as the special case where $\eta_{t}$ is constant, which we denote as CHNG. The HNG option pricing model of Heston and Nandi, (2000) is also nested by imposing $\eta_{t}=1$. We refer to our new model as DHNG model, where “D” stands for dynamic, and highlight its key properties below.
@@
We derive a closed-form CBOE VIX pricing formula for the DHNG model and develop an analytical option pricing formula, which constitutes a separate and important methodological contribution. When volatility risk aversion follows a stochastic process, the conditional moment-generating function (MGF) of future cumulative returns does not have an affine form, making it very challenging to derive option prices. However, we introduce a novel approximation method that is applicable to non-affine models. This approach constructs an auxiliary MGF for a simplified, related problem and then applies a Taylor expansion to obtain an analytical option pricing formula. The approximation method is shown to be highly accurate in an empirically relevant simulation design.
@@
To implement the theoretical framework, we connect the model for $\eta_{t}$ with observed data using an observation-driven approach, which is a natural choice since the GARCH model under the physical measure also follows an observation-driven formulation. Specifically, we propose a score-driven model following Creal et al., (2013), where the first-order conditions of the log-likelihood function define the innovations in the dynamic model for $\eta_{t}$. This design is intuitive because $\eta_{t}$ is updated to minimize pricing errors. Conveniently, this structure also enhances robustness to model misspecification. Another advantage of the observation-driven model is its simplicity in estimation, which is particularly useful in applications involving long sample periods and large panels of option prices.
@@
An empirical application using 32 years of daily S&P 500 returns, the CBOE VIX, and a large panel of option prices provides strong support for the model. The estimation incorporates information from both the physical and risk-neutral measures, as advocated by Chernov and Ghysels, (2000). The results show that the pricing kernel with time-varying volatility risk aversion typically reduces pricing errors by 50% or more, as measured by the root mean square error. This reduction is achieved for both VIX and option prices, in both in-sample and out-of-sample comparisons. The estimated time variation in the shape of the pricing kernel aligns with previous studies, generally exhibiting a U-shape, but taking on an inverted U-shape during certain periods, including 2004–2007 and around 1993 and 2017.
@@
Finally, having established the VRR as a fundamental quantity, it is natural to investigate its connection to key asset pricing variables, including those linked to pricing kernel puzzles. Theoretical and empirical studies suggest that heterogeneous beliefs and disagreements about the physical distribution can contribute to a U-shaped pricing kernel, see e.g. Shefrin, (2001, 2008), Bakshi and Madan, (2008), Bakshi et al., (2010). Similar arguments have been made about sentiment and uncertainty among investors, see e.g. Han, (2008), Polkovnichenko and Zhao, (2013), Baker et al., (2016), and Bollerslev et al., (2018). Bali and Zhou, (2016) argue that market uncertainty can be approximated by the variance risk premium, see Carr and Wu, (2008). The VRR is obviously related to the variance risk premium, which has been shown to predict stock returns, see Bollerslev et al., (2009). A straightforward explanation is that preferences are state-dependent, with market uncertainty being a possible state variable, see e.g. Grith et al., (2017). We connect our results to these studies by showing that the monthly average VRR is closely related to commonly used measures of sentiment, uncertainty, and disagreement, as well as other key economic indicators.
Our paper is related to and builds on a large body of literature, including studies that have sought to address pricing kernel puzzles by augmenting existing models with additional state variables, see e.g. Chabi-Yo et al., (2008), Chabi-Yo, (2012), Brown and Jackwerth, (2012) and Song and Xiu, (2016). In our framework, the quantity VRR can be viewed as an additional state variable that emerges naturally from the model structure.
@@
Our work is also related to Barone-Adesi et al., (2008), who estimated a GJR-GARCH model for returns under the physical measure. They did not specify a pricing kernel but simply assumed that the model under the risk-neutral measure is also a GJR-GARCH model, with time-varying parameters calibrated to minimize option pricing errors. While we also seek to minimize pricing errors, our approach is fundamentally different. Our starting point is a pricing kernel and a model for the physical probability measure, $\mathbb{P}$, and the two imply the model under the risk-neutral probability measure, $\mathbb{Q}$. An empirical comparison in Christoffersen et al., (2013) explores the same idea as in Barone-Adesi et al., (2008), as they estimate separate HNG models under both $\mathbb{P}$ and $\mathbb{Q}$ measures. They referred to this structure as an ad-hoc model, because it is not derived from their pricing kernel. However, we show that the pricing kernel with a time-varying VRR is incoherent with HNG models (with constant parameters) under both $\mathbb{P}$ and $\mathbb{Q}$ measures.
@@
Additionally, our paper contributes to the literature on time-varying risk aversion, which relates to the local shape of the pricing kernel. A seminal paper by Campbell and Cochrane, (1999) addresses this concept, while Li, (2007) demonstrated how time-varying risk aversion links to a time-varying risk premium. Moreover, González et al., (2018) identified risk aversion dynamics as a key determinant of stock market betas. More recently, Bekaert et al., (2020) introduced a no-arbitrage asset pricing model with time-varying risk aversion for pricing equities and corporate bonds, where risk aversion is driven by uncertainty shocks.
@@
The rest of the paper is organized as follows. We present the theoretical model in Section 1 and derive closed-form pricing formula for the VIX and an option pricing formula, based on the new approximation method in Section 2. Section 3 introduces a dynamic model for the variance risk ratio using a score-driven approach. An empirical application with 32 years of S&P 500 returns, VIX, and option prices is presented in Section 4. We show that the variance risk ratio relates to well-known measures of sentiment, disagreement, and uncertainty in Section 5. A summary is presented in Section 6. All proofs are provided in the Appendix A and supplementary empirical results can be found in the Online Appendix.
@@ 1 The Model
The observed variables include the daily returns, $R_{t}\equiv\log\left(S_{t}/S_{t-1}\right)$, where $S_{t}$ is the underlying asset price, and a vector of derivative prices, $X_{t}$. We will work with two filtrations, $\mathcal{F}_{t}=\sigma(\{R_{j},X_{j}\},j\leq t)$ and $\mathcal{G}_{t}=\sigma(\{R_{j},X_{j-1}\},j\leq t)$, where the latter arises naturally from a factorization of the joint likelihood function. Clearly, $\mathcal{F}_{t-1}\subset\mathcal{G}_{t}\subset\mathcal{F}_{t}$ for all $t$. Throughout this paper, we will use the notation $\mathbb{E}_{t}^{\mathbb{P}}\left(\cdot\right)\equiv\mathbb{E}^{\mathbb{P}}\left(\cdot|\mathcal{G}_{t}\right)$ and $\mathbb{E}_{t}^{\mathbb{Q}}\left(\cdot\right)\equiv\mathbb{E}_{t}^{\mathbb{Q}}\left(\cdot|\mathcal{G}_{t}\right)$ to denote the conditional expectations with respect to $\mathcal{G}_{t}$ under $\mathbb{P}$ and $\mathbb{Q}$ measures, respectively.
We model returns in physical measure $\mathbb{P}$ using the classical Heston-Nandi GARCH model (HNG), which has a convenient structure for derivatives pricing. This model is given by
@@
where the return shock, $z_{t+1}$, is independent and identically distributed (iid) with a standard normal distribution, $N(0,1)$; $r$ is the risk-free rate; $\lambda$ is the equity risk premium because the expected return is given by $\mathbb{E}^{\mathbb{P}}\left(\exp\left(R_{t+1}\right)|\mathcal{F}_{t}\right)=\exp\left(r+\lambda h_{t+1}\right)$; and $h_{t+1}=\mathrm{var}^{\mathbb{P}}(R_{t+1}|\mathcal{F}_{t})$ is the daily conditional variance, see Heston and Nandi, (2000). The HNG model captures time variation in the conditional variance, as is the case for ARCH and GARCH models, see Engle, (1982) and Bollerslev, (1986). However, the HNG model also allows for dependence between returns and volatility (leverage). It is straightforward to verify that $\mathrm{cov}^{\mathbb{P}}(R_{t},h_{t+1}|\mathcal{F}_{t-1})=-2\gamma\alpha h_{t}$, which shows that the magnitude of the leverage effect is defined by $\gamma$. A leverage effect is required to generate the empirically important volatility smirk in option prices. The dynamic structure of the HNG model is carefully crafted to yield a closed-form solution for option valuation, and Heston and Nandi, (2000) showed that the continuous limit of the HNG model (as the time interval between observations shrinks to zero) yields a variance process, $h_{t}$, that converges weakly to a continuous-time square-root variance process, see Feller, (1951), Cox et al., (1985), and Heston, (1993).